OpenStudy (anonymous):

I am on FLVS doing some testing and this question is killing me, since I am not feeling good, I need heeeeeeeelp. The question is. Cheyanne is saving for a car. The amount she has saved for the car, f(x), can be represented by the function f(x) = 65x + 750, where x is the number of weeks she has saved. Suppose 750 changed to 700. What does this mean in the context of the problem?

OpenStudy (campbell_st):

well 750 is the y intercept for the linear equation... so in week 0 she has $750 in her bank account... so if this amount is reduced to $700 what effect will it have on her savings... that's what you are being asked

OpenStudy (anonymous):

do you have answer choices

OpenStudy (anonymous):

Since the 750 in the equation is the initial amount she saved, when 750 becomes 700, it means the initial amount decreased from $750 to $700. For the second question: The point slope form of the equation of a line is y−y1=m(x−x1) where m = slope, and the point is (x1,y1) Now use the equation above, and just substitute the given values.

OpenStudy (anonymous):

it should be Changing 750 to 700 means the initial amount she saved decreased.